# Restoring Missing Modes of 21cm Intensity Mapping with Deep Learning: Impact on BAO Reconstruction

### v1

@@ 3.3 BAO Reconstruction
  • Calculate the density map $\rho_{i}$ from distribution of particles, using Cloud-in-Cloud algorithm (CIC).
@@
Since the product of a 21cm intensity mapping experiment is a temperature map, it is necessary to convert the temperature fluctuation into distribution of particles. To achieve this, we employ supersampling of the density field to generate particles. To maintain the high resolution in the particle representation and ensure the smoothness of the field, the number of particles, $N_{\rm part}$, is set to be at least $2^{3}$ times greater than the number of pixels in the original field. Furthermore, since the reconstruction algorithm does not differentiate between the dark matter field and any biased tracers, it is crucial to de-bias the temperature distribution so that the input map provided to the algorithm accurately represents an estimation of the underlying dark matter distribution. For such purpose, we define the input density distribution
$\hat{\delta}_{\rm input}(\mathbf{x})=\delta T(\mathbf{x})/b_{T}.$ (3.8)
@@
Here, $b_{T}$ is the linear bias coefficient, estimated from the cross-correlation between dark matter field and the temperature fluctuation map. Specifically, we averaged the scale-dependent bias $b_{T}(k)=\langle\delta_{\rm dm}\delta T\rangle/\langle\delta_{\rm dm}\delta_{\rm dm}\rangle$ up to a cut-off scale $k_{\rm bcut}=0.035~{}{\rm h/Mpc}$. In practice, without direct access to the dark matter distribution, the bias parameter must be measured using alternative modeling methods, which can introduce additional errors. As discussed in [71], an inaccurate estimation of the bias parameter may result in further damping of the BAO signature. We will explore these observational effects in future discussions.
@@ 4 RESULT
Due to GPU constraints, the AI model was trained at a low resolution of $N_{\rm grid}=128^{3}$. However, this low resolution could significantly impact the performance of BAO reconstruction, which is our main objective. Fortunately, we discovered an very interesting scale invariance in the AI model when applied to the LSS data. Specifically, we found that the model, despite being trained on low resolution data, could be directly applied to higher resolution maps as long as the dimension of the dimension of the high resolution map can be processed by the network architecture depicted in Figure 1. In Figure 2, we present a visual comparison of true temperature maps (top panels), observed maps with foreground-contaminated modes removed (middle panels), and the AI restored map (lower panels). Different columns display the same map at varying grid sizes, ranging from $64^{3}$, to $128^{3}$, $256^{3}$, and $320^{3}$. As demonstrated, in all these cases, the AI-restored maps successfully resemble the true brightness temperature maps in each column.
Before delving into the detailed performance analysis in each case, let us first understand the reasons behind such scale invariance of AI model. This property can largely be attributed to two factors: the self-similarity exhibited by large-scale structure evolution and the hierarchical structure of U-Net model. Specifically, from the perspective of perturbation theory, the mode-coupling coefficients appeared in the non-linear dynamical evolution of the perturbation [72]
@@
Here $P_{12}(\mathbf{k})=\langle\delta_{1}\delta_{2}\rangle(\mathbf{k})$ represents the cross power spectrum between two fields, and $P_{1/2}(\mathbf{k})=\langle\delta_{1/2}\delta_{1/2}\rangle(\mathbf{k})$ is the respective auto power spectra. Figure 4 illustrates the cross-correlation ratio in two-dimensional $k_{\perp}-k_{\parallel}$ space, with resolutions varying from $64^{3}$ on the left to $320^{3}$ on the right. As shown, the cross-correlation ratio for the low-resolution $64^{3}$ sample is approximately $\sim 0.8$ at very large scales ($k\lesssim 0.1~{}{\rm h/Mpc}$) but drops to around $\sim 0.5$ at $k\sim 0.2-0.3~{}{\rm h/Mpc}$. Additionally, because the map is trained in the configuration space, the $r(\mathbf{k})$ for modes unaffected by the foreground in the upper left corner does not perfectly equal one. For the grid size of $128^{3}$, the correlation ratio with the true temperature map significantly improves, reaching around $0.8$ at $k\lesssim 0.3~{}{\rm h/Mpc}$. With larger grid sizes, the performance further improves. For example, the $256^{3}$ grid map achieves a correlation ratio of approximately $0.8$ at $k\sim~{}0.6{\rm h/Mpc}$, compared to about $0.5-0.6$ for the $128^{3}$ grid map. This trend continues with a finer $320^{3}$ grid, which is the maximum size our GPU memory can accommodate. Therefore, it is reasonable to expect that, with more advanced GPU capabilities, the model could potentially achieve even better results.
Furthermore, we also present the one-dimensional correlation ratio $r(k)$, averaging over results in Figure 4. As displayed in Figure 5, solid lines represent the correlation ratio for the AI-restored fields, while dashed lines indicate the ratio for maps with missing modes due to foreground avoidance. Due to the extension of the foreground wedge to larger $k$ values, the 1D correlation ratio does not converge to one, even at smaller scales. Except for the lowest resolution map of $64^{3}$, the overall correlation ratio of the AI-restored fields reaches approximately $\sim 0.9$ at very small scales, $k\sim 1~{}{\rm h/Mpc}$. This performance exceeds the scale limits ($\sim 0.6~{}{\rm h/Mpc}$) where the BAO reconstruction algorithms typically cease to be effective, demonstrating that our AI restoration maintains sufficient fidelity for subsequent application of BAO reconstruction.
Figure 6: The BAO signature for various fields. The upper panel illustrates the BAO signal for the true brightness temperature and the AI-restored $T_{b}$ both before and after BAO reconstruction, with all signals here derived from data with a grid size of $320^{3}$. Comparing the post-BAO reconstructed signals (thick black solid line vs. red dotted line), it is evident that the AI restoration has minimal impact on the reconstruction quality. The lower panel highlights the differences between the AI-restored BAO signal and the true signal, both before (black solid line) and after BAO reconstruction (purple solid line).
@@ 4.1 BAO Reconstruction
Finally, having restored the lost Fourier modes, we proceed to reconstruct the linear BAO signature using our algorithm. To more clearly observe the effects on the BAO, we apply our trained AI model as well as the subsequent BAO reconstruction to a pair of simulations that share the same initial condition seed but differ in their transfer functions: one retains the BAO signature, while the other employs a ‘no-wiggle’ smooth function. This allows us to define the BAO signature as
$S_{\rm BAO}(k)=\frac{P_{\rm BAO}-P_{\rm nwig}}{P_{\rm nwig}}(k)$ (4.3)
in Fourier space, where $P_{\rm BAO}$ is the power spectrum with BAO, and $P_{\rm nwig}$ is the ‘no-wiggle’ power spectrum.
The results are displayed in the upper panel of Figure 6. All signals, except for the linear model (black dotted line), are derived from data with a grid size of $320^{3}$. As seen, the BAO signal from the AI-restored field (thin green solid line) is very similar to the true 21cm brightness temperature map (thin orange solid line), indicating the effectiveness of our AI model. Due to the non-linear evolution, the wiggles after the third peaks are dampened away compared to the linear model. the AI-restored field (red dotted line) closely aligns with the BAO-reconstructed true brightness temperature that includes all modes (thick black solid line). This suggests that our AI restoration has minimal impact on the effectiveness of BAO reconstruction. In both fields, BAO reconstruction significantly enhances the BAO signature, especially beyond the third peak, where it had been dampened in the original field. In the lower panel of the figure, we present the differences in the BAO signature between the AI-restored and true temperature fields, both before (black solid line) and after BAO reconstruction (purple solid line). Notably, the difference in both scenarios is around $1\%$, which is relatively small. Additionally, at larger scales ($k\lesssim 0.2{\rm h/Mpc}$), the two curves largely overlap, indicating that the BAO reconstruction does not introduce significant numerical errors. However, at smaller scales, although the error range remains similar, the curves begin to diverge, suggesting that non-linear BAO reconstruction introduces its own numerical errors in the field and BAO signatures.
@@
One might notice that both reconstructed BAO signatures deviate from the linear theory (black dotted line), particularly at the second and third peaks. This deviation can be attributed primarily to the effects of clustering bias. Despite correcting for the linear bias factor before reconstruction, the relationship between the brightness temperature and the underlying dark matter is complex, involving various non-linear and non-local contributions. As previous studies have shown [39], such complexities can diminish the effectiveness of the reconstruction. A secondary factor may be the resolution of the field. Constrained by the GPU limitation, grid size of $320^{3}$ is not ideal for the BAO reconstruction. However, such effect is less significant compared to the clustering bias. Moreover, future advancements in computational capabilities are expected to mitigate the influence of resolution constraints on subsequent BAO reconstructions.

### v2

@@ 3.3 BAO Reconstruction
  • Calculate the density map $\rho_{i}$ from the distribution of particles, using the Cloud-in-Cloud algorithm (CIC).
@@
Since the product of a 21cm intensity mapping experiment is a temperature map, it is necessary to convert the temperature fluctuation into distribution of particles. To achieve this, we employ supersampling of the density field to generate particles. To maintain high resolution in the particle representation and ensure the smoothness of the field, the number of particles, $N_{\rm part}$, is set to be at least $2^{3}$ times greater than the number of pixels in the original field. Furthermore, since the reconstruction algorithm does not differentiate between the dark matter field and any biased tracers, it is crucial to de-bias the temperature distribution so that the input map provided to the algorithm accurately represents an estimation of the underlying dark matter distribution. For such purpose, we define the input density distribution
$\hat{\delta}_{\mathrm{input}}(\mathbf{x})=\delta T(\mathbf{x})/\left(\bar{T}_{\mathrm{HI}}b_{\mathrm{T}}\right).$ (3.8)
@@
Here, $\bar{T}_{\mathrm{HI}}$ is the average HI temperature, $b_{T}$ is the linear bias coefficient, estimated from the cross-correlation between dark matter field and the temperature fluctuation map. Specifically, we averaged the scale-dependent $\bar{T}_{\mathrm{HI}}b_{\mathrm{T}}(k)=\langle\delta_{\mathrm{dm}}\delta T\rangle/\langle\delta_{\mathrm{dm}}\delta_{\mathrm{dm}}\rangle$ up to a cut-off scale $k_{\mathrm{bcut}}=0.035~{}{\rm h/Mpc}$. In practice, without direct access to the dark matter distribution, the factor $\bar{T}_{\mathrm{HI}}b_{\mathrm{T}}$ must be measured using alternative modeling methods, which can introduce additional errors. Fortunately, as proposed by [41], one can iteratively reconstruct the field with different bias factors (or the combined term $\bar{T}_{\mathrm{HI}}b_{\mathrm{T}}$ here) and minimize BAO smearing by simultaneously using both the peak location and width of the BAO signature. However, this approach is beyond the scope of the current paper and will be explored in future work.
Figure 2: Visual comparison between the true brightness temperature (upper panels), the observed temperature with intrinsic and foreground wedge removed (middle panels), and the U-Net restored temperature (lower panels). Different columns demonstrate maps at various resolutions, ranging from $64^{3}$ (first column) to $128^{3}$ (second), $256^{3}$ (third), and $320^{3}$ (last column). Here, our low resolution results ($64^{3}$ and $128^{3}$) were obtained using the stage-one AI model trained at a resolution of $128^{3}$, while the higher-resolution results ($256^{3}$ and $320^{3}$) were produced using the stage-two model, which was retrained with $256^{3}$ data. (The same applies to the results in the following.) All brightness temperatures are in unit of ${\rm mK}$.
Figure 3: Comparison of the two-dimensional power spectrum $P(k_{\perp},k_{\parallel})$ of 21cm brightness temperature across various fields and resolutions. The first row displays the power spectrum of the true brightness temperature. The second row presents the observed map, with modes contaminated by foregrounds removed, specifically those with $k_{\parallel}<0.1~{}{\rm h/Mpc}$ and within the wedge. The last row displays the power restored by the U-Net model. Additionally, different columns correspond to power spectra calculated from maps at varying resolutions, starting from $64^{3}$ up to $320^{3}$.
3.4 BAO Extraction and Parametric Fitting
After reconstruction, we examine the BAO signature of the processed power spectrum to assess the reconstruction performance. A commonly used approach is to normalize by the power spectrum of seed-matched no-wiggle simulations, allowing the BAO signature to be defined as
$S^{\mathrm{sim}}_{\mathrm{BAO}}(k)=\frac{P_{\mathrm{BAO}}-P_{\mathrm{nwig}}}{P_{\mathrm{nwig}}}(k),$ (3.9)
where $P_{\rm BAO}$ represents the power spectrum with BAO, and $P_{\rm nwig}$ denotes the corresponding ‘no-wiggle’ power spectrum, and the superscript $\mathrm{sim}$ indicates that this BAO signal is derived directly from simulation-based manipulation. To examine the performance of our AI model under the most idealized conditions, we first process both $P_{\mathrm{BAO}}$ and its corresponding $P_{\mathrm{nwig}}$ through our AI restoration and BAO reconstruction pipeline. This allows us to establish a baseline for evaluating the performance of the combined AI restoration and BAO reconstruction procedure.
Of course, in real observations, one does not have access to a Universe with a no-wiggle power spectrum. Since various effects can distort the overall shape of the processed power spectrum, evaluating the performance of BAO reconstruction becomes challenging. Therefore, alternative methods are required to robustly extract the BAO signal. Consequently, we adopt the standard BAO measurement approach [74], fitting the observationally processed power spectrum with a template that accounts for the BAO signature, the overall broadband shape, and additional distortions introduced by processing pipeline. Specifically, we model the template power spectrum as
$P^{\mathrm{temp}}(k)=P^{\mathrm{temp}}_{\mathrm{nwig}}(k)T(k)\left[S^{\mathrm{temp}}_{\mathrm{BAO}}(k)+1\right].$ (3.10)
Here, $P^{\mathrm{temp}}_{\mathrm{nwig}}(k)$ is the no-wiggle power spectrum template, unaffected by foreground removal or AI processing, $T(k)$ is a fourth-order polynomial
$T(k)=\sum_{i=0}^{4}a_{i}k^{i},$ (3.11)
accounting for broadband shape distortion. And finally, $S^{\mathrm{temp}}_{\mathrm{BAO}}(k)$ is the theoretical BAO template defined as
$S^{\mathrm{temp}}_{\mathrm{BAO}}(k)=S^{\mathrm{temp}}_{\mathrm{BAO,lin}}(k)\exp\left[-k^{2}\Sigma^{2}/2\right].$ (3.12)
Here, $S^{\mathrm{temp}}_{\mathrm{BAO,lin}}(k)$ is the theoretical linear BAO signal obtained by dividing the output of the Boltzmann code CAMB [75] by a smooth, fitted no-wiggle power spectrum. The parameter $\Sigma$ quantifies the degree of nonlinear smearing or the effect of reconstruction. Therefore, compared with Equation (3.9), this approach follow a more observationally motivated procedure and avoids relying on AI- and BAO-reconstructed no-wiggle simulations. All AI-induced shape distortions in the band power are absorbed into the transfer function $T(k)$, ensuring an unbiased evaluation of the process. Meanwhile, the effectiveness of the BAO reconstruction—with or without AI restoration—is reflected in the fitted damping parameter $\Sigma$. As we will show, improved BAO reconstruction results in a smaller $\Sigma$, indicating reduced nonlinear smearing.
Notice that since we are focusing on evaluating the AI restoration in the context of BAO reconstruction, we do not include cosmology dependent scaling factors that account for the geometric shift of the wavenumber $k$. Additionally, for simplicity, we use only the diagonal components of the covariance matrix in the $\chi^{2}$ estimation, which was computed using our original $37$ simulated training samples.
@@ 4 RESULT
4.1 AI-restored Maps
In Figure 2, we present a visual comparison of true temperature maps (top panels), observed maps with foreground-contaminated modes removed (middle panels), and the AI-restored map (lower panels). Different columns display the same map at varying grid sizes, ranging from $64^{3}$, to $128^{3}$, $256^{3}$, and $320^{3}$. Due to a slight performance drop of the stage-two retrained AI model on lower-resolution maps, the $64^{3}$ and $128^{3}$ results were obtained using the stage-one AI model trained at $128^{3}$, while the higher-resolution results ($256^{3}$ and $320^{3}$) were generated using the stage-two model retrained with $256^{3}$ data. As demonstrated, in all these cases, the AI-restored maps successfully resemble the true brightness temperature maps in each column.
Despite variations in performance when applying our AI model to maps of different resolutions, it is noteworthy that it consistently succeeds in recovering the missing modes. Before delving into the detailed performance analysis in each case, let us first explore the underlying reasons behind this apparent scale invariance of the AI model. This property can largely be attributed to two factors: the self-similarity exhibited by the evolution of large-scale structure and the hierarchical structure of the U-Net model. Specifically, from the perspective of perturbation theory, the mode-coupling coefficients appeared in the non-linear dynamical evolution of the perturbation [76]
@@ 4.1 AI-restored Maps
Here $P_{12}(\mathbf{k})=\langle\delta_{1}\delta_{2}\rangle(\mathbf{k})$ represents the cross power spectrum between two fields, and $P_{1/2}(\mathbf{k})=\langle\delta_{1/2}\delta_{1/2}\rangle(\mathbf{k})$ is the respective auto power spectra. Figure 4 illustrates the cross-correlation ratio in two-dimensional $k_{\perp}-k_{\parallel}$ space, with resolutions varying from $64^{3}$ on the left to $320^{3}$ on the right. As shown, the cross-correlation ratio for the low-resolution $64^{3}$ sample is approximately $\sim 0.8$ at very large scales ($k\lesssim 0.1~{}{\rm h/Mpc}$) but drops to around $\sim 0.5$ at $k\sim 0.2-0.3~{}{\rm h/Mpc}$. Additionally, because the map is trained in the configuration space, the $r(\mathbf{k})$ for modes unaffected by the foreground in the upper left corner does not perfectly equal one. For the grid size of $128^{3}$, the correlation ratio with the true temperature map significantly improves, reaching around $0.8$ at $k\lesssim 0.3~{}{\rm h/Mpc}$. With larger grid sizes, the performance further improves. For example, the $256^{3}$ grid map achieves a correlation ratio of approximately $0.8$ at $k\sim 0.6~{}{\rm h/Mpc}$, compared to about $0.6$ for the $128^{3}$ grid map. This trend continues with a finer $320^{3}$ grid, which is the maximum size our GPU memory can accommodate. Therefore, it is reasonable to expect that, with more advanced GPU capabilities, the model could potentially achieve even better results.
Furthermore, we also present the one-dimensional correlation ratio $r(k)$, averaging over the results in Figure 4. As displayed in Figure 5, solid lines represent the correlation ratio for the AI-restored fields, while dashed lines indicate the ratio for maps with missing modes due to foreground avoidance. Due to the extension of the foreground wedge to larger $k$ values, the 1D correlation ratio does not converge to one, even at smaller scales. Except for the lowest resolution map of $64^{3}$, the overall correlation ratio of the AI-restored fields reaches approximately $\sim 0.9$ at very small scales, $k\sim 1~{}{\rm h/Mpc}$. This performance exceeds the scale limits ($\sim 0.6~{}{\rm h/Mpc}$) where the BAO reconstruction algorithms typically cease to be effective, demonstrating that our AI restoration maintains sufficient fidelity for subsequent application of BAO reconstruction.
@@ 4.2 BAO Reconstruction
Having restored the lost Fourier modes, we proceed to reconstruct the BAO signature using our algorithm. As discussed in Section 3.4, we present the BAO result in two different ways. First, as a baseline evaluation of the most idealized scenario, we apply our trained AI model as well as the subsequent BAO reconstruction to a pair of simulations that share the same initial condition seed but differ in their transfer functions: one retains the BAO signature, while the other employs a ‘no-wiggle’ smooth function. The results are displayed in panel (a) of Figure 6. All signals, except for the linear model (thin black dotted line), are derived from data with a grid size of $256^{3}$. As seen, the BAO signal from the AI-restored field (thick sky-blue solid line) is very similar to the true 21cm brightness temperature map (thin blue dash-dotted line), indicating the effectiveness of our AI model. Due to the non-linear evolution, the wiggles after the third peaks are dampened away compared to the linear model. Similarly, after BAO reconstruction, we find that the AI-restored field (red solid line) closely aligns with the true post-reconstruction BAO signal that includes all modes (thick black dotted line). This suggests that, in this most idealized scenario, AI restoration does not appear to significantly impact the overall effectiveness of BAO reconstruction. As shown, the BAO reconstruction enhances the BAO signature, particularly beyond the third peak, where it was previously dampened in the original field.
Figure 6: The BAO signatures of various fields. From panel (a) to (c), the BAO signature is obtained by dividing the power spectrum with BAO features by a seed-matched no-wiggle power spectrum, with each panel corresponding to a different level of processing applied to the no-wiggle spectrum. Panel (a): The baseline result where the $P_{\mathrm{nwig}}$ has undergone the same AI restoration and/or BAO reconstruction process as the corresponding power spectrum with the BAO signature. As shown, in this most idealized case, AI restoration does not appear to significantly affect the performance of BAO reconstruction. Panel (b): A more realistic scenario where the no-wiggle spectrum has not undergone AI restoration. In this case, the AI restoration introduces mild shape deviations to the BAO signature. Panel (c): A scenario where the no-wiggle spectrum has undergone neither AI restoration nor BAO reconstruction. Now, the BAO reconstruction introduce quite significant shape deviation. Panel (d): BAO signals obtained from MCMC fitting of Equation (3.12), with the smearing parameter $\Sigma$ determined from the fit. Here, the template $P_{\mathrm{nwig}}$ has not undergone any AI restoration or BAO reconstruction, and the BAO template $S_{\mathrm{BAO}}$ is defined using the theoretical linear model.
Of course, above result does not imply that the sharpened BAO peak can be extracted without difficulty. Since both the AI model and BAO reconstruction introduce scale-dependent distortions in the recovered power spectrum, additional care is required. In a gradually more realistic scenario, shown in the second row of Figure 6, we present the BAO signature obtained by dividing the processed power spectrum by a no-wiggle spectrum that has undergone no AI restoration (panel b), and one that has undergone neither AI restoration nor BAO reconstruction (panel c). In panel (b), the sky-blue solid line is calculated by dividing the AI-restored power spectrum by a no-wiggle $P(k)$ without AI restoration. It shows that AI restoration alone introduces a mild shape distortion, leading to slight modifications in the BAO signature. These distortions are carried through to the BAO reconstruction, as seen in the red solid line, which is obtained by dividing the processed power spectrum by a no-wiggle spectrum that has undergone BAO reconstruction but not AI restoration. Nevertheless, it is evident that the BAO reconstruction still effectively reduces nonlinear smearing, allowing more BAO peaks to be recovered. In panel (c), we further present the BAO signature obtained by dividing the processed power spectrum by a no-wiggle $P(k)$ that has undergone neither AI restoration nor BAO reconstruction. In this case, both lines are divided by the same unprocessed no-wiggle spectrum, which also serves as the no-wiggle template $P^{\mathrm{temp}}_{\mathrm{nwig}}(k)$ described in Equation (3.10). As a result, the sky-blue solid line in this panel is identical to the corresponding line in panel (b). However, the red solid line, representing the post-BAO-reconstruction result, shows that BAO reconstruction along introduces significantly greater shape deviations compared to AI restoration. Addressing these shape distortions introduced by BAO reconstruction is beyond the scope of this paper, but we plan to investigate them further and explore possible improvements in future work.
To isolate the BAO signature from the distorted curves above, we follow the fitting procedure described in Section 3.4 and extract the smearing parameter $\Sigma$. As previously mentioned, the template $P_{\mathrm{nwig}}^{\mathrm{temp}}(k)$ has not undergone any AI restoration or BAO reconstruction, and the BAO template $S^{\mathrm{temp}}_{\mathrm{BAO}}(k)$ is computed using the output of the Boltzmann code CAMB along with the corresponding fitted no-wiggle power spectrum. As a result, the fitted smearing parameter $\Sigma$ serves as a proxy for the amplitude of the BAO peaks, providing a quantitative measure of the (in)effectiveness of the BAO reconstruction. Specifically, we employed the Markov Chain Monte Carlo (MCMC) method to fit the corresponding power spectrum. The extracted BAO signatures are presented in panel (d) of Figure 6, with each curve showing the fitted BAO, as described by Equation (3.12), and the corresponding $\Sigma$ value from the best fit to the simulated data. Compared to panel (a), this approach accurately captures the smearing parameter $\Sigma$, and the resulting BAO signal closely matches the corresponding result shown in panel (a). Additionally, in Appendix A, we also present corner plots of all parameters for two specific scenarios: the AI-restored $P(k)$ without BAO reconstruction (Figure 7) and with BAO reconstruction (Figure 8). As shown, The posterior distribution of the polynomial coefficients $a_{i}$ is quite Gaussian, whereas the smearing coefficient exhibits some non-Gaussianity. Due to the large error on the power spectrum arising from the size of our simulation box, the one-$sigma$ error on $\Sigma$ is quite large. Nevertheless, as shown in Figure 6, the best-fit BAO model remains sufficiently accurate.
@@
One might notice that both reconstructed BAO signatures deviate from the linear theory (thin black dotted line), particularly at the second and third peaks. This deviation primarily arises from the imperfect de-biasing process (Equation 3.8). Although the linear factor $\bar{T}_{\mathrm{HI}}b_{\mathrm{T}}$ has already been corrected for, the relationship between the brightness temperature fluctuation and the underlying dark matter is quite complex, various non-linear and non-local contributions may still reduce the effectiveness of the reconstruction. A secondary factor may be the resolution of the field. Constrained by the GPU limitation, the current grid size is not ideal for the BAO reconstruction. However, such effect is less significant compared to the clustering bias. Moreover, future advancements in computational capabilities are expected to mitigate the influence of resolution constraints on subsequent BAO reconstructions.